OpenStudy (anonymous):

how can I explain using z^n=r^n(cos(ntheta)+isin(theta)) that z^1/n=r^1/n(cos((ntheta+2kpi)/n)+isin((theta+2kpi)/n)))? and that k= 0,1,2,...,n-1 (so we get n solutions out)

OpenStudy (camper4834):

whoah we are going to need to clean this up a little

OpenStudy (camper4834):

\[z^n = r^{n} \cos(n \theta)+i \sin(\theta)\] \[\frac{ z^1 }{ n } = \frac{ r^1 }{ n }\cos(\frac{ n \theta + 2k \pi }{ n })+ i \sin(\frac{ \theta+2k \pi }{ n })\]

OpenStudy (camper4834):

did i do that right?

OpenStudy (anonymous):

yes! besides the ntheta after the cos. it should be just theta

OpenStudy (camper4834):

for both equations?

OpenStudy (anonymous):

no just the bottom. because it becomes 1/n instead of n

OpenStudy (camper4834):

\[z^n = r^n(\cos(n \theta)+i \sin(\theta))\] but this is correct?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

wait no. there is n in the sin

OpenStudy (camper4834):

\[\frac{ z^1 }{ n }=\frac{ r^1 }{ n }(\cos(\frac{ \theta + 2k \pi }{ n })+i \sin(\frac{ \theta +2k \pi }{ n }))\]

OpenStudy (anonymous):

exactly

OpenStudy (camper4834):

\[z^n=r^n(\cos(n \theta)+i \sin(n \theta))\]

OpenStudy (camper4834):

okay

OpenStudy (anonymous):

so then i need to use the top equation to show how you get bottom one (where its 1/n)

OpenStudy (camper4834):

\[z^n=r^n(\cos(n \theta)+i \sin(n \theta))\] \[\frac{ z^1 }{ n }=\frac{ r^1 }{ n }(\cos(\frac{ \theta + 2k \pi }{ n })+i \sin(\frac{ \theta +2k \pi }{ n }))\] k = 0, 1, 2 ... n-1

OpenStudy (camper4834):

im assuming it is an identity

ganeshie8 (ganeshie8):

it is used for taking nth root of a complex number

OpenStudy (camper4834):

where z could be 1+3i?

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

but how do I show that one leads to the other one? like where does the nkpi comes from? (i understand that it's adding another circle but how do I explain it in the equation? and why do I get n-1 solutions from it

ganeshie8 (ganeshie8):

first of all, you're adding 2kpi to the numearator inside angle, not nkpi

OpenStudy (anonymous):

oops i meant 2kpi

ganeshie8 (ganeshie8):

lets start from De Moivre's theorem for integer powers : \[\large z^n = r^{n} \cos(n \theta)+i \sin(n\theta)\]

ganeshie8 (ganeshie8):

is that what you want to start with ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

if a function is periodic with a period of \(T\), can we say : \[f(x) = f(x+T)\] cuz its value repeats every \(T\) units of x

OpenStudy (anonymous):

2pi isnt it?

ganeshie8 (ganeshie8):

yes period of \(\cos (\theta)\) is \(2\pi\) so can we write \(\cos \theta = \cos (\theta + 2\pi)\) ?

ganeshie8 (ganeshie8):

and how about : \[\cos \theta = \cos (\theta + 2\pi) = \cos(\theta +2\pi + 2\pi+2\pi + \cdots)\] ?

OpenStudy (anonymous):

yea i get this part

ganeshie8 (ganeshie8):

instead of that messy looking expression, we say : \[\cos(\theta) = \cos(\theta + 2k\pi)\] where k = 0,1,2,3,... (also negatives)

ganeshie8 (ganeshie8):

good, lets start with a complex number

ganeshie8 (ganeshie8):

\[z = \cos(\theta) + i\sin(\theta)\] which is same as \[z = \cos(\theta + 2k\pi) + i\sin(\theta + 2k\pi)\] yes ?

ganeshie8 (ganeshie8):

because both sin and cos are periodic and ofcourse k is some integer

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

*any integer

ganeshie8 (ganeshie8):

*all integers

ganeshie8 (ganeshie8):

good, next rise it to 1/n power using De moivre's thm

ganeshie8 (ganeshie8):

\[z = \cos(\theta) + i\sin(\theta)\] which is same as \[z = \cos(\theta + 2k\pi) + i\sin(\theta + 2k\pi)\] By De Moivre's thm : \[z = \cos(\theta) + i\sin(\theta)\] which is same as \[z^{1/n} = \cos\left(\dfrac{\theta + 2k\pi}{n}\right) + i\sin\left(\dfrac{\theta + 2k\pi}{n}\right)\]

ganeshie8 (ganeshie8):

k can be ANY integer, however any "n" consecutive values of "k" gives u distinct solutions

ganeshie8 (ganeshie8):

"n" consecutive values of k could be : k = 0,1,2,3,... n-1 or k= 1,2,3,4,...n or any other sequence of "n" consecutive integers will do

OpenStudy (anonymous):

this is the part I don't understand how come you get n solutions for every equation

ganeshie8 (ganeshie8):

thats a good question, lets see the solutions in complex plane

OpenStudy (anonymous):

lets say if I had n=4, i will plug in k as 0,1,2,3 and get four solutions.

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

yes and those 4 solutions will be DISTINCT meaning, they will be 4 different points in complex plane

ganeshie8 (ganeshie8):

lets take a number and take its nth root

ganeshie8 (ganeshie8):

pick some number

OpenStudy (anonymous):

why